# Duality for Fitting’s Multi-valued Modal logic via bitopology and bi-Vietoris coalgebra

### v2

@@ 3.2 Functors
We claim that $\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle=\langle\Box T_{1}(a)\rightarrow 0\rangle$. Let $\mathcal{W}\in\langle\Box T_{1}(a)\rightarrow 0\rangle$. Then $\mathcal{W}(\Box T_{1}(a)\rightarrow 0)=1$. Hence, $\mathcal{W}(\Box T_{1}(a))=0$. Using the Kripke condition, we have, $0=\mathcal{W}(\Box T_{1}(a))=\bigwedge\{\mathcal{U}(T_{1}(a)):\mathcal{W}\mathcal{R}_{\Box}\mathcal{U}\}$. Since $\mathcal{U}(T_{1}(a))=0\text{ or }1$, hence $\exists$ $\mathcal{U}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ with $\mathcal{W}\mathcal{R}_{\Box}\mathcal{U}$ such that $\mathcal{U}(T_{1}(a))=0$. Then $\mathcal{U}\in\langle T_{1}(a)\rightarrow 0\rangle$. Therefore, $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle T_{1}(a)\rightarrow 0\rangle\neq\emptyset$. Thus, $\mathcal{W}\in\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle$. Similarly, by employing the Kripke condition, we can show that if $\mathcal{W}\in\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle$ then $\mathcal{W}\in\langle\Box T_{1}(a)\rightarrow 0\rangle$. Since $\langle\Box T_{1}(a)\rightarrow 0\rangle\in\beta_{2}$, we have $\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle\in\beta_{2}$. As a result, $[\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$.
@@
$\psi(T_{\ell^{*}}(a^{*})\rightarrow a^{*})=\left\{\begin{array}[]{ll}\ell^{*}&\text{ if }\psi(a^{*})=\ell^{*}\\ 1&\text{ if }\psi(a^{*})\neq\ell^{*}\\ \end{array}\right.$
@@ 3.3 Bitopological Duality for $\mathcal{L}$$-\mathcal{ML}$ Algebras
Define $\zeta_{(P,\alpha_{P},\mathcal{R})}:(P,\alpha_{P},\mathcal{R})\to\mathcal{G}\circ\mathcal{F}(P,\alpha_{P},\mathcal{R})$ by $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\psi)=\psi(p)$, where $p\in P$ and $\psi\in HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Theorem 1 shows that $\zeta_{(P,\alpha_{P},\mathcal{R})}$ is a bi-homeomorphism in the category $PBS_{\mathcal{L}}$. We show that $\zeta_{(P,\alpha_{P},\mathcal{R})}$ and $\zeta^{-1}_{(P,\alpha_{P},\mathcal{R})}$ satisfy the conditions given in item 2 of Definition 10. We claim that for any $p,p^{\prime}\in P$, $p^{\prime}\in\mathcal{R}[p]\iff\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. Let $p^{\prime}\in\mathcal{R}[p]$. Suppose $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}\psi)\geq\ell$, where $\ell\in\mathcal{L}$ and $\psi\in HOM_{{PBS}_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Then $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}\psi)=(\Box_{\mathcal{R}}\psi)(p)=\bigwedge\{\psi(p^{*}):p*\in\mathcal{R}[p]\}$. Since $p^{\prime}\in\mathcal{R}[p]$ and $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}\psi)\geq\ell$, we have $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})(\psi)\geq\ell$. Hence, $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)\mathcal{R}_{\Box_{\mathcal{R}}}\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})$, i.e., $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. Now we verify if $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$ then $p^{\prime}\in\mathcal{R}[p]$. We verify its contrapositive statement. Suppose $p^{\prime}\notin\mathcal{R}[p]$. By Definition 10, $\mathcal{R}[p]$ is a pairwise compact subset of pairwise Boolean space $P$. Then It is straightforward to demonstrate that $\mathcal{R}[p]$ is pairwise closed. Therefore we can get a $\tau_{1}^{P}$-basis open set $\mathcal{O}\in\beta_{1}^{P}$ such that $p^{\prime}\in\mathcal{O}$ and $\mathcal{O}\subseteq P-\mathcal{R}[p]$, i.e., $\mathcal{O}\cap\mathcal{R}[p]=\emptyset$. Define a mapping $f:P\to\mathcal{L}$ by
$f(p)=\left\{\begin{array}[]{ll}0&\text{ if $p\in\mathcal{O}$}\\ 1&\text{ if $p\in\mathcal{O}^{c}$}\\ \end{array}\right.$
Then $f$ is a pairwise continuous map from $(P,\tau_{1}^{P},\tau_{2}^{P})$ to $(\mathcal{L},\tau,\tau)$. As a result, it can be shown that $f\in HOM_{PBS_{\mathcal{L}}}((P,\alpha_{P}),(\mathcal{L},\alpha_{\mathcal{L}}))$. Now, $\Box_{\mathcal{R}}f(p)=\bigwedge\{f(z):z\in\mathcal{R}[p]\}=1$ and $f(p^{\prime})=0$. Hence, $\zeta_{(P,\alpha_{P},\mathcal{R})}(p)(\Box_{\mathcal{R}}f)=1$ but $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})(f)\neq 1$. Therefore, $\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\notin\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. Hence, we have for any $p,p^{\prime}\in P$, $p^{\prime}\in\mathcal{R}[p]\iff\zeta_{(P,\alpha_{P},\mathcal{R})}(p^{\prime})\in\mathcal{R}_{\Box_{\mathcal{R}}}[\zeta_{(P,\alpha_{P},\mathcal{R})}(p)]$. As a result, $\zeta_{(P,\alpha_{P},\mathcal{R})}$ and $\zeta^{-1}_{(P,\alpha_{P},\mathcal{R})}$ satisfy conditions $(i)$ and $(ii)$ mentioned in item 2 of Definition 10. ∎
@@ 4 Coalgebraic Duality Theorem for Fitting’s multi-valued logic
This section intends to define an endofunctor called as $\mathcal{L}$-biVietoris functor, $V_{\mathcal{L}}^{bi}:PBS_{\mathcal{L}}\to PBS_{\mathcal{L}}$. After that, we exhibit that the category $PRBS_{\mathcal{L}}$ is isomorphic to the category $COALG(V_{\mathcal{L}}^{bi})$ of coalgebras for the endofunctor $V_{\mathcal{L}}^{bi}$.
@@ 4.1 The structure of the endofunctor $V_{\mathcal{L}}^{bi}$
We define the pairwise Vietoris space as follows:
@@
Let $(S,\tau_{1}^{S},\tau_{2}^{S})$ be a pairwise topological space and $\mathcal{K}(S)$ the set of all pairwise closed subsets of $S$. We define $\Box U=\{C\in\mathcal{K}(S):C\subseteq U\}$ and $\Diamond U=\{C\in\mathcal{K}(S):C\cap U\neq\emptyset\}$, $U\subseteq S$. Let $\beta_{1}^{S}$ and $\beta_{2}^{S}$ be the basis for the topologies $\tau_{1}^{S}$ and $\tau_{2}^{S}$, respectively. The pairwise Vietoris space $V_{P}(S)$ of the pairwise topological space $(S,\tau_{1}^{S},\tau_{2}^{S})$ is defined as a pairwise topological space $(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$, where $\tau_{1}^{V}$ is the topology on $\mathcal{K}(S)$ generated by subbasis $\{\Box U,\Diamond U:U\in\beta_{1}^{S}\}$ and the topology $\tau_{2}^{V}$ on $\mathcal{K}(S)$ is generated by subbasis $\{\Box U,\Diamond U:U\in\beta_{2}^{S}\}$.
We then show that $V_{P}(S)$ is a pairwise Boolean space whenever $S$ is a pairwise Boolean space.
@@
Proof.
We shall show that $\beta_{1}^{V}=\tau_{1}^{V}\cap\delta_{2}^{V}$ is a basis for $\tau_{1}^{V}$, where $\delta_{2}^{V}$ is the set of $\tau_{2}^{V}$-closed sets. Let $\mathcal{O}\in\tau_{1}^{V}$. Then $\mathcal{O}$ can be expressed as $\mathcal{O}=\displaystyle\bigcup_{\lambda\in\Lambda}(\bigcap_{r=1}^{n_{\lambda}}\Box C_{r}\cap\bigcap_{t=1}^{m_{\lambda}}\Diamond C_{t})$, $C_{r},C_{t}\in\beta_{1}^{S}=\tau_{1}^{S}\cap\delta_{2}^{S}$. To demonstrate that $\beta_{1}^{V}$ is a basis for $\tau_{1}^{V}$, we have to prove that $\bigcap_{r=1}^{n_{\lambda}}\Box C_{r}\cap\bigcap_{t=1}^{m_{\lambda}}\Diamond C_{t}\in\beta_{1}^{V}$. Because the finite intersection of the members of $\beta_{1}^{V}$ is again in $\beta_{1}^{V}$, it is sufficient to establish that for $C\in\beta_{1}^{S}$, $\Box C,\Diamond C\in\beta_{1}^{V}$. As $\tau_{1}^{V}$ is the topology generated by the subbasis $\{\Box C,\Diamond C:C\in\beta_{1}^{S}\}$, hence $\Box C,\Diamond C\in\tau_{1}^{V}$. Now we see that $(\Box C)^{c}=\Diamond C^{c}$ and $(\Diamond C)^{c}=\Box C^{c}$. Since $C\in\beta_{1}^{S}$, so $C^{c}\in\beta_{2}^{S}$. As a result, $\Box C,\Diamond C\in\delta_{2}^{V}$. Henceforth, $\Box C,\Diamond C\in\beta_{1}^{V}$. Similarly, it can be shown that $\beta_{2}^{V}=\tau_{2}^{V}\cap\delta_{1}^{V}$, $\delta_{1}^{V}$ is the set of $\tau_{1}^{V}$-closed sets, is a basis for $\tau_{2}^{V}$. ∎
@@
Proof.
Let $C,C^{\prime}\in\mathcal{K}(S)$ and $C\neq C^{\prime}$. Let $z\in C$ such that $z\neq z^{\prime}$, $\forall z^{\prime}\in C^{\prime}$. For each point $z^{\prime}\in C^{\prime}$, we choose disjoint open sets $U_{z^{\prime}}^{c}\in\beta_{2}^{S}$ and $U_{z^{\prime}}\in\beta_{1}^{S}$ (using the condition that $(S,\tau_{1}^{S},\tau_{2}^{S})$ is pairwise Hausdorff space.) containing points $z^{\prime}$ and $z$, respectively. So the collection $\{U_{z^{\prime}}^{c}:z^{\prime}\in C^{\prime}\}$ is $\tau_{2}^{S}$-open covering of $C^{\prime}$. As $C^{\prime}$ is pairwise compact, so there is a finite collection $\{U_{z^{\prime}_{i}}^{c}:i=1,2,\cdots,n\}$ such that $C^{\prime}\subseteq\bigcup_{i=1}^{n}U_{z^{\prime}_{i}}^{c}$. Let $V^{\prime}=\displaystyle\bigcup_{i=1}^{n}U_{z^{\prime}_{i}}^{c}$ and $U=\displaystyle\bigcap_{i=1}^{n}U_{z^{\prime}_{i}}$. As $z\in C\cap U$, hence $C\cap U\neq\emptyset$. Also, $C^{\prime}\cap U=\emptyset$ because $C^{\prime}\subseteq U^{c}$. It follows that $C\in\Diamond U\in\tau_{1}^{V}$ and $C^{\prime}\notin\Diamond U$ i.e., $C^{\prime}\in(\Diamond U)^{c}=\Box U^{c}\in\tau_{2}^{V}$. So we have two disjoint open sets $\Diamond U\in\tau_{1}^{V}$ and $\Box U^{c}\in\tau_{2}^{V}$ containing $C$ and $C^{\prime}$, respectively. ∎
@@
Proof.
It is known from Proposition 1 that $\{\Box U,\Diamond U:U\in\beta_{1}^{S}\cup\beta_{2}^{S}\}$ is a subbasis for the topology $\tau_{1}^{S}\vee\tau_{2}^{S}$. We shall show that every cover of $\mathcal{K}(S)$ by subbasis-open sets has a finite subcover. Let $\mathcal{K}(S)=\bigcup_{\lambda\in\Lambda}\Box U_{\lambda}\cup\bigcup_{i\in I}\Diamond V_{i}$. Consider $S_{1}=S-\bigcup_{i\in I}V_{i}$. Then $S_{1}$ is a pairwise closed subset of $S$. Hence, $S_{1}\in\mathcal{K}(S)$. Since, $S_{1}\notin\Diamond V_{i}$ for each $i\in I$, so that $S_{1}\in\bigcup_{\lambda\in\Lambda}\Box U_{\lambda}$. Then for some $\lambda^{\prime}\in\Lambda$, $S_{1}\in\Box U_{{\lambda}^{\prime}}$. As a result, $S_{1}\subseteq U_{{\lambda}^{\prime}}$ and hence $S-U_{{\lambda}^{\prime}}\subseteq S-S_{1}=\displaystyle\bigcup_{i\in I}V_{i}$. Then, $S=U_{{\lambda}^{\prime}}\cup\displaystyle\bigcup_{i\in I}V_{i}$. As $S$ is pairwise compact, we have $S=U_{\lambda^{\prime}}\cup\bigcup_{i=1}^{n}V_{i}$. Let $A$ be an arbitrary element of $\mathcal{K}(S)$. If $A\subseteq U_{{\lambda}^{\prime}}$ then $A\in\Box U_{{\lambda}^{\prime}}$ otherwise $A\subseteq\bigcup_{i\in I}V_{i}$ i.e., $A\cap V_{i}\neq\emptyset$ for some $i\in\{1,2,\cdots,n\}$. As a result, $A\in\Box U_{\lambda^{\prime}}\cup\bigcup_{i\in I}\Diamond V_{i}$. Therefore, $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is pairwise compact. ∎
@@
If $(S,\tau_{1}^{S},\tau_{2}^{S})$ is a pairwise Boolean space then $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is also a pairwise Boolean space.
We now construct the $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}$.
@@
Given that $f$ is a pairwise continuous map from a pairwise Boolean space $S_{1}$ to a pairwise Boolean space $S_{2}$. Let $K\in V_{P}(S_{1})$. Then $K$ is a pairwise closed subset of $S_{1}$ and hence $K$ is pairwise compact. Now $V_{\mathcal{L}}^{bi}(\eta)(K)=\eta[K]$ is a pairwise compact subset of $S_{2}$. Since $S_{2}$ is a pairwise Boolean space, $\eta[K]$ is a pairwise closed subset of $S_{2}$. As a result, $V_{\mathcal{L}}^{bi}(\eta)(K)\in V_{P}(S_{2})$. To show that $V_{\mathcal{L}}^{bi}(\eta)$ is pairwise continuous, let $C\in\beta_{1}^{S_{2}}$ and $D\in\beta_{2}^{S_{2}}$. Then
@@
Let $(S,\alpha_{S},\mathcal{R})$ be an object in $PRBS_{\mathcal{L}}$. Then $\mathfrak{B}(S,\alpha_{S},\mathcal{R})$ is an object in $COALG(V_{\mathcal{L}}^{bi})$
@@
We shall show that $\mathcal{R}[-]:(S,\alpha_{S})\to V_{\mathcal{L}}^{bi}(S,\alpha_{S})$ is an arrow in $PBS_{\mathcal{L}}$. By the conditions given in the object part of Definition 10, we know that for each $s\in S$, $\mathcal{R}[s]$ is pairwise compact subset of $S$. As $S$ is pairwise Boolean space, hence $\mathcal{R}[s]$ is a pairwise closed subset of $S$. Thus $\mathcal{R}[s]\in V_{P}(S)$. Let $U\in\beta_{1}^{S}$. Then
@@
Finally, based on the preceding theorems, we can conclude:
@@
Fitting’s multi-valued modal logic is sound and complete in the sense of coalgebras of the biVietoris functor $V_{\mathcal{L}}^{bi}$.
@@
and $0,1$ are considered as constants i.e., nullary-operations. Algebras of Łukasiewicz-$n$-valued modal logic can be found in [13].

### v3

@@ 3.2 Functors
We claim that $\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle=\langle\Box T_{1}(a)\rightarrow 0\rangle$. Let $\mathcal{W}\in\langle\Box T_{1}(a)\rightarrow 0\rangle$. Then $\mathcal{W}(\Box T_{1}(a)\rightarrow 0)=1$. Hence, $\mathcal{W}(\Box T_{1}(a))=0$. Using the Kripke condition, we have, $0=\mathcal{W}(\Box T_{1}(a))=\bigwedge\{\mathcal{U}(T_{1}(a)):\mathcal{W}\mathcal{R}_{\Box}\mathcal{U}\}$. Since $\mathcal{U}(T_{1}(a))=0\text{ or }1$, hence $\exists$ $\mathcal{U}\in HOM_{\mathcal{VA}_{\mathcal{L}}}(\mathcal{A},\mathcal{L})$ with $\mathcal{W}\mathcal{R}_{\Box}\mathcal{U}$ such that $\mathcal{U}(T_{1}(a))=0$. Then $\mathcal{U}\in\langle T_{1}(a)\rightarrow 0\rangle$. Therefore, $\mathcal{R}_{\Box}[\mathcal{W}]\cap\langle T_{1}(a)\rightarrow 0\rangle\neq\emptyset$. Thus, $\mathcal{W}\in\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle$. Similarly, by employing the Kripke condition, we can show that if $\mathcal{W}\in\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle$ then $\mathcal{W}\in\langle\Box T_{1}(a)\rightarrow 0\rangle$. Since $\langle\Box T_{1}(a)\rightarrow 0\rangle\in\beta_{2}$, we have $\langle\mathcal{R}_{\Box}\rangle\langle T_{1}(a)\rightarrow 0\rangle\in\beta_{2}$. As a result, $\mathcal{R}_{\Box}]\langle a\rangle\in\beta_{1}$.
@@
[…]
@@ 3.3 Bitopological Duality for $\mathcal{L}$-$\mathcal{ML}$ Algebras
[…]
@@ 4 Coalgebraic Duality Theorem for Fitting’s multi-valued logic
The aim of this section is to adapt the bi-Vietoris construction of Lauridsen on pairwise Stone spaces [7] to our $\mathcal{L}$-valued pairwise Boolean spaces. We thereby define an endofunctor $V_{\mathcal{L}}^{bi}$ on the category $PBS_{\mathcal{L}}$ by incorporating the structure map, and we show that $V_{\mathcal{L}}^{bi}$ preserves the $\mathcal{L}$-valued structure. After that, we exhibit that the category $PRBS_{\mathcal{L}}$ is isomorphic to the category $COALG(V_{\mathcal{L}}^{bi})$ of coalgebras for the endofunctor $V_{\mathcal{L}}^{bi}$, and combine this identity with the bitopological duality developed in Section 3 to establish a coalgebraic duality $\mathcal{MA}_{\mathcal{L}}\simeq COALG(V_{\mathcal{L}}^{bi})^{op}$ from which soundness and completeness of Fitting’s $\mathcal{L}$-valued modal logic with respect to $V_{\mathcal{L}}^{bi}$-coalgebras follow.
In this framework, our assumption over $\mathcal{L}$, the algebra of truth values, is same as in previuos sections. We work with Fitting’s-style $\mathcal{L}$-valued modal logic that is modified by Maruyama in his work [12]. The underlying algebra of the $\mathcal{L}$-$\mathcal{VL}$-algebra is Heyting, and hence pseudo-complement exists. However, the modal signature in this framework is negation-free, i.e., neither Boolean nor involutive negation. The modal operator $\Box$ preserves finite meets and $\top$ element, and commutes with each $U_{\ell}$, $\ell\in\mathcal{L}$.
@@ 4.1 The structure of the endofunctor $V_{\mathcal{L}}^{bi}$
To set up the coalgebraic presentation of our bitopological duality for Fitting’s $\mathcal{L}$-valued modal logic, we recall the pairwise Vietoris space construction and its basic properties on pairwise Stone spaces from [7].
The notion of pairwise Vietoris space is given in the following definition:
@@
[7] Let $(S,\tau_{1}^{S},\tau_{2}^{S})$ be a pairwise topological space and $\mathcal{K}(S)$ the set of all pairwise closed subsets of $S$. We define $\Box U=\{C\in\mathcal{K}(S):C\subseteq U\}$ and $\Diamond U=\{C\in\mathcal{K}(S):C\cap U\neq\emptyset\}$, $U\subseteq S$. Let $\beta_{1}^{S}$ and $\beta_{2}^{S}$ be the basis for the topologies $\tau_{1}^{S}$ and $\tau_{2}^{S}$, respectively. The pairwise Vietoris space $V_{P}(S)$ of the pairwise topological space $(S,\tau_{1}^{S},\tau_{2}^{S})$ is defined as a pairwise topological space $(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$, where $\tau_{1}^{V}$ is the topology on $\mathcal{K}(S)$ generated by subbasis $\{\Box U,\Diamond U:U\in\beta_{1}^{S}\}$ and the topology $\tau_{2}^{V}$ on $\mathcal{K}(S)$ is generated by subbasis $\{\Box U,\Diamond U:U\in\beta_{2}^{S}\}$.
Remark 1.
In the above definition of $V_{P}(S)$, $\Box U$ and $\Diamond U$ denote the subbasic open sets. The algebras of Fitting’s $\mathcal{L}$-valued modal logic have only one modal operator $\Box$, and its semantics is given via subbasic open set $\Box U$ . We do not use $\Diamond$ as the algebraic signature. Thus, $\Diamond$ acts as a topological generator and doesn’t match up with any algebraic modality in our work.
The following three lemmas (zero-dimensionality, Hausdorffness, compactness of $V_{P}(S)$, whenever $S$ is a pairwise Boolean space) are of routine in the pairwise Vietoris set-up and can be found in Lauridsen’s work [7]. Together they yield that $V_{P}(S)$ is a pairwise Boolean space whenever $S$ is a pairwise Boolean space.
@@
[7] If $(S,\tau_{1}^{S},\tau_{2}^{S})$ is a pairwise Boolean space then $V_{P}(S)=(\mathcal{K}(S),\tau_{1}^{V},\tau_{2}^{V})$ is also a pairwise Boolean space.
We now extend the construction of bi-Vietoris functor on pairwise Stone spaces (cf. [7]) to our $\mathcal{L}$-valued pairwise Boolean spaces setting by integrating the subalgebra-indexed, meet-preserving structure map $\alpha$ through $V_{P}$. Consequently, we obtain the $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}$ on $PBS_{\mathcal{L}}$ and use it to prove the coalgebraic duality for the algebras of Fitting’s $\mathcal{L}$-valued modal logic.
@@
Remark 2.
We note that for $\mathcal{L}=\{0,1\}$, the two-element Boolean algebra, our $\mathcal{L}$-biVietoris structure recovers Lauridsen’s bi-Vietoris construction, and if both the topologies $\tau_{1}$ and $\tau_{2}$ are equal then $\mathcal{L}$-biVietoris functor becomes classical Vietoris functor on the category of Stone spaces.
@@
Given that $\eta$ is a pairwise continuous map from a pairwise Boolean space $S_{1}$ to a pairwise Boolean space $S_{2}$. Let $K\in V_{P}(S_{1})$. Then $K$ is a pairwise closed subset of $S_{1}$ and hence $K$ is pairwise compact. Now $V_{\mathcal{L}}^{bi}(\eta)(K)=\eta[K]$ is a pairwise compact subset of $S_{2}$. Since $S_{2}$ is a pairwise Boolean space, $\eta[K]$ is a pairwise closed subset of $S_{2}$. As a result, $V_{\mathcal{L}}^{bi}(\eta)(K)\in V_{P}(S_{2})$. To show that $V_{\mathcal{L}}^{bi}(\eta)$ is pairwise continuous, let $C\in\beta_{1}^{S_{2}}$ and $D\in\beta_{2}^{S_{2}}$. Then
@@
Let $(S,\alpha_{S},\mathcal{R})$ be an object in $PRBS_{\mathcal{L}}$. Then $\mathfrak{B}(S,\alpha_{S},\mathcal{R})$ is an object in $COALG(V_{\mathcal{L}}^{bi})$ .
@@
We shall show that $\mathcal{R}[-]:(S,\alpha_{S})\to V_{\mathcal{L}}^{bi}(S,\alpha_{S})$ is an arrow in $PBS_{\mathcal{L}}$. By the conditions given in the object part of Definition 10, we know that for each $\in S, [s]\mathcal{R}[s$] is pairwise compact subset of $S$. As $S$ is pairwise Boolean space, hence $\mathcal{R}[s]$ is a pairwise closed subset of $S$. Thus $\mathcal{R}[s]\in V_{P}(S)$. Let $U\in\beta_{1}^{S}$. Then
@@
For $\mathcal{L}=\mathbf{2}$, the two-element Boolean algebra, our duality which is established in Section 3 gives the Jónsson-Tarski duality for modal algebras. When composed with Theorem 7, this provides $\mathcal{MA}_{\mathbf{2}}\simeq COALG(V_{\mathbf{2}}^{bi})^{op}$, which recovers the Abramsky-Kupke-Kurz-Venema coalgebraic duality.
Finally, based on the preceding theorems (Theorems 4,6,7), we can conclude:
@@
Fitting’s multi-valued modal logic is sound and complete in the sense of coalgebras of the $\mathcal{L}$-biVietoris functor $V_{\mathcal{L}}^{bi}$.
@@
and $0,1$ are considered as constants, i.e., nullary-operations. Algebras of Łukasiewicz-$n$-valued modal logic can be found in [14].